# Out of Style: RAG’s Fragility to Linguistic Variation

### v1

@@ 1 Introduction
Retrieval-augmented Generation (RAG) systems enhance Large Language Models (LLMs) by integrating external knowledge retrieval, grounding their output in factual context to improve accuracy and reliability (Lewis et al., 2021; Gao et al., 2024). However, their widespread integration into real-world applications (K2view, 2024) introduce potential challenges regarding robustness to linguistic variations. Users bring varied backgrounds, domains, and cultural contexts that naturally produce linguistic differences in their queries (Park et al., 2024; Li et al., 2020; Lorenzo-Dus & Bou-Franch, 2013). As Figure 1 illustrates, different from the carefully curated queries from traditional NLP benchmarks, real-world user-LLM queries tend to be less formal and frequently contain grammatical inconsistencies (Ouyang et al., 2023). Failing to account for these linguistic variations risks excluding a broad segment of users from effective interaction, especially for users whose linguistic expressions fall outside the narrow patterns these systems are tuned on (Liang et al., 2023). Moreover, unlike standalone LLMs, RAG systems incorporate multiple interdependent components, making them susceptible to cascading errors arising at both the retrieval and generation stages (Asai et al., 2023; Yoran et al., 2024a; Kim et al., 2025). A truly robust RAG system should maintain consistent retrieval effectiveness and generation quality across the full spectrum of user linguistic variations.
@@
We present a large-scale systematic investigation of how variations in linguistic characteristics affect the robustness of RAG systems. We focus on formality, readability, politeness, and grammatical correctness - fundamental aspects of linguistic variations that are critical and common in real-world user-LLM queries (Li et al., 2024; Yin et al., 2024; Hagen et al., 2024). By automatically rewriting queries across these dimensions, we analyze how linguistic variations impact each RAG system component, as well as potential cascading errors throughout the pipeline. Our evaluation encompasses two retrieval models, namely Contriever (Izacard et al., 2021) and ModernBERT (Nussbaum et al., 2024), and nine LLMs from three families (Llama 3.1, Qwen 2.5, Gemma 2) of varying scales across four open-domain Question Answering (QA) datasets: PopQA (Mallen et al., 2023), EntityQuestions (Sciavolino et al., 2022), MS MARCO (Bajaj et al., 2018), and Natural Questions (Kwiatkowski et al., 2019a).
@@
Furthermore, RAG systems exhibit greater vulnerability to linguistic variations than LLM-only generations, suggesting cascading errors across components. The PopQA dataset shows a 22.53% average performance drop in RAG systems versus only 10.78% in LLM-only generations. These findings highlight the urgent need to improve the robustness of retrieval components when handling linguistically varied queries. We also find that while advanced RAG techniques such as query expansion with HyDE (Gao et al., 2022) and documents reranking tend to improve overall performance, they remain similarly vulnerable to linguistic variations.
@@ 2 Related Work
Prior research investigated retrieval system robustness to noisy queries (Campos et al., 2023; Chen et al., 2022; 2023b) and specific linguistic variations including word substitutions (Wu et al., 2022), aspect changes and paraphrasing (Penha et al., 2022), typographical errors (Zhuang & Zuccon, 2021), and grammatical variations (Long et al., 2024). Our work provides the first holistic evaluation of RAG systems’ robustness to diverse linguistic variations and uncovers cascading failures across the entire RAG pipeline.
@@
Prior research has examined impacts of syntactic perturbations Moradi & Samwald (2021); Singh et al. (2024), round-trip translation Bhandari & Chen (2023), politeness variations Yin et al. (2024), equivalent queries Cao et al. (2024) and scale of model size on robustness to grammatical errors (Yin et al., 2020; Hagen et al., 2024). Rawte et al. (2023) examined how formality and readability affect LLM performance in isolation. In contrast, our research investigates a broader spectrum of linguistic variations, generates data for each variation using LLM-based rewrites, and studies their compounding effects throughout the end-to-end RAG pipeline comprehensively.
@@
While RAG systems demonstrate impressive performance (Lewis et al., 2021; Gao et al., 2024) and reduced hallucinations (Mallen et al., 2023), vulnerabilities exist with increasing retrieval context noise (Chen et al., 2023a), irrelevant contexts (Yoran et al., 2024b), and noise impacts (Fang et al., 2024). Yang et al. (2025) analyzed how spurious features affect RAG perform. These studies focus primarily on retrieved content noise rather than initial query variations. Our work uniquely demonstrates how diverse linguistic variations in user queries compound throughout the RAG pipeline, exposing critical vulnerabilities in systems serving diverse users.
@@ 3 Robustness Evaluation Approach
Given a seed dataset $\mathcal{D}=\{(x_{1},y_{1}),(x_{2},y_{2}),\ldots\}$, where $x_{i}$ and $y_{i}$ indicate $i$-th input and output, respectively, we reformulate each query $x_{i}\rightarrow x_{i}^{\prime}$ based on four linguistic aspects. A robust RAG system, composed of a retriever $\mathcal{R}$ and a generator $\mathcal{G}$ operating on corpus $\mathcal{C}$, should maintain performance when processing linguistically varied inputs. For the retrieval component, we expect retrieved documents $\mathbf{D}_{i}=\mathcal{R}(x_{i},\mathcal{C})$ and $\mathbf{D}_{i}^{\prime}=\mathcal{R}(x_{i}^{\prime},\mathcal{C})$ to both contain the answer $y_{i}$. For generation, given retrieved documents $\mathbf{D}_{i},\mathbf{D}_{i}^{\prime}$, a robust system should produce: $\mathcal{G}(x_{i},\mathbf{D}_{i})\approx\mathcal{G}(x_{i}^{\prime},\mathbf{D}_{i}^{\prime})\approx y_{i}$.
@@ 3.1 Linguistic Variations
Formality in language lacks universal definition (Pavlick & Tetreault, 2016; Mosquera & Moreda, 2011; Fang & Cao, 2009), but encompasses situational factors (Hovy, 1987; Lahiri et al., 2011), grammar quality (Peterson et al., 2011), and specific linguistic elements like contractions (Heylighen et al., 1999). We quantify formality using the RoBERTa-based formality classifier from (Babakov et al., 2023b).
@@
Readability quantifies text comprehensibility through complexity and linguistic features. We employ the widely-used Flesch Reading Ease Score (FRES; Flesch 1948) to assess readability (Rawte et al., 2023; Han et al., 2024), defined in A.1.
@@
In our work, we define grammatical correctness as the preservation of both grammaticality (Chomsky, 2002) and semantic fidelity. We alter the grammatical correctness through two approaches, inspired by Yin et al. (2020); Zhuang & Zuccon (2021); Lichtarge et al. (2019): (1) Typos, where random addition, deletion, or substitution operations are applied at a 20% probability per word, requiring an edit distance of at least 1; and (2) Round-trip translation (RTT) via English-Afrikaans-English using EasyNMT with opus-mt model, requiring the output to not be the exact same as the original query.
@@ 3.2 Query Rewriting
We systematically reformulate queries $x_{i}\rightarrow x_{i}^{\prime}$ across targeted linguistic dimensions while ensuring all rewrites satisfy dimension-specific thresholds and preserve semantic meaning (as specified in Appendix A.1). For each distinct dimension—formality, readability, and politeness—we take $5,000$ original queries and use GPT-4o-mini (OpenAI et al., 2024) to generate rewrites using three different prompts (detailed in Appendix G). We design distinct query sets across all datasets to be less formal, less readable, or more polite—directions chosen because the well-curated original datasets contained predominantly formal, readable, and neutral language, allowing our modifications to explore more realistic linguistic variations. This creates $15,000$ rewritten samples per dimension for each dataset, over which we average results to account for prompt stochasticity. For grammatical correctness, we similarly use $5,000$ original queries but apply deterministic transformations through typo introduction and round-trip translation as described in Section 3.1, creating $5,000$ modified samples for each approach. We ensure that grammatically modified queries differ from originals using GLEU and edit distance measurements. Example rewrites are shown in Table 1.
For each individual linguistic dimension and dataset, we construct controlled comparative datasets $\mathcal{D}^{\prime}=\{(x_{1}^{\prime},y_{1}),(x_{2}^{\prime},y_{2}),\ldots\}$. To validate our approach, we annotate $100$ queries across all linguistic variations, finding that 97.33% of the rewritten queries preserve the semantic meaning of the original query. Details are available in Appendix A.2.
@@
Grammar $\downarrow$: Round-Trip Translation Original: Which company is HMS Blankney produced by?
@@
Grammar $\downarrow$: Typos Original: when did the japanese river otter become extinct
@@ 4.1 Benchmarks
In this work, we use four open-domain QA datasets as seed datasets $\mathcal{D}$ and evaluate the effects how linguistic variations of the original queries affect RAG systems. PopQA (Mallen et al., 2023) is a large-scale entity-centric open-domain QA dataset about entities with a wide variety of popularity. EntityQuestions (Sciavolino et al., 2022) is a set of simple, entity-rich questions based on facts of Wikidata. MS MARCO (Bajaj et al., 2018) contains questions derived from real user search queries from Bing’s search logs. Natural Questions (Kwiatkowski et al., 2019a) contains questions consisting of real, anonymized, aggregated queries to the Google search engine. Both PopQA and EntityQuestions consist of well-structured, standardized, and simple queries, while queries from MS MARCO and Natural Questions exhibit free-form and arbitrary. We evaluate using PopQA’s test split, EntityQuestions’ dev split, and both dev and test splits from MS MARCO and Natural Questions.
@@ 4.2 Models
We evaluate nine most advanced open-source LLMs from three model families with various scales: Llama 3.1 (Llama-3.1-8B-Instruct, Llama-3.1-70B-InstructGrattafiori et al. 2024), Qwen 2.5 (Qwen2.5-3B-Instruct, Qwen2.5-7B-Instruct, Qwen2.5-32B-Instruct, Qwen2.5-72B-Instruct;  Yang et al. 2024; Team 2024), and Gemma 2 (gemma-2-2b-it, gemma-2-9b-it, gemma-2-27b-itTeam et al. 2024). Detailed hyperparameter settings are included in the Appendix B.1.
We use few-shot prompting to ensure that the model outputs are in the correct format. For each dataset and linguistic characteristic, we include two question-answer pairs in the context: one random original query with its answer, and its corresponding linguistically rewritten version with the same answer. This balanced approach exposes the model to both original and rewritten query formats to ensure fairness. We also include the top five retrieved passages in the context. Detailed prompts are provided in Appendix H.
@@ 4.3 Metrics
The generation stage is assessed using a comprehensive set of metrics: Answer Match (AM) measures the percentage of the predictions that any substring of the prediction is an exact match of any of the ground truth answers, Exact Match (EM) measures the percentage of predictions that exactly match any of the ground truth answers, and F1 Score (F1) captures the harmonic mean of precision and recall in generated responses.
@@ 5.1 Retrieval Analysis
Our analysis (Table 2) reveals significant linguistic fragility in retrieval systems, with performance degradation averaging 16.7% for Contriever and 13.3% for ModernBERT across all modifications. Results show highest sensitivity on PopQA (19.78% average impact), particularly to grammatical transformations (29.34% from RTT). MS MARCO exhibits the second-highest impact (16.78%), with striking sensitivity to formality changes (40.41% with Contriever), suggesting retrieval systems may be implicitly optimized for specific linguistic patterns, limiting effectiveness when handling diverse query variations.
@@
Retrieval performance drops independent of query length.
Our analysis, as shown in Table 2 demonstrates that query length changes do not directly correlate with retrieval performance. Queries with increased formality showed substantial length increases (+13.65 tokens) yet produced inconsistent performance impacts across datasets. Conversely, round-trip translated queries were marginally shorter (-0.26 tokens) but consistently caused significant performance degradation. This indicates retrieval models respond more to linguistic quality (grammatical correctness, readability) than to query length itself, highlighting the need for systems robust to linguistic variations rather than optimized for specific query lengths.
Scaling up number of documents improves performance.
Table 2 shows performance degradation ($\Delta$R@K) decreasing as $K$ increases, indicating linguistic perturbations cause relevant documents to slide down rather than disappear from the ranked list. As an example, for rewrites with typos, $\Delta$ R@K for Contriever decreases from 22.2 at R@5 to 12.0 at R@100, and for ModernBERT from 20.1 to 9.8. This is detailed further in Appendix C.1. This ranking deterioration forces downstream language models to operate with suboptimal information, potentially compromising response quality. We further investigate this hypothesis in Section 5.4.2 by examining if rerankers can improve recall scores for Top-5 retrieved documents.
5.2 Generation Analysis
The RAG experiment results on ModernBERT retriever and nine LLMs are presented in Table 3. Overall, RAG systems show performance degradation on all linguistic variations.

### v2

@@ Abstract
Out of Style: RAG’s Fragility to Linguistic Variation
Tianyu Cao1thanks: Equal contribution.   Neel Bhandari1∗   Akhila Yerukola1   Akari Asai12   Maarten Sap1 1 Language Technologies Institute, Carnegie Mellon University 2 Allen Institute for AI {tianyuca, neelbhan, ayerukol, aasai, msap2}@cs.cmu.edu
@@ 1 Introduction
Retrieval-augmented Generation (RAG) systems enhance Large Language Models (LLMs) by integrating external knowledge retrieval, grounding their output in factual context to improve accuracy and reliability (Lewis et al., 2021; Gao et al., 2024). However, their widespread integration into real-world applications (K2view, 2024) introduces potential challenges regarding robustness to linguistic variations. Users bring varied backgrounds, domains, and cultural contexts that naturally produce linguistic differences in their queries (Park et al., 2024; Li et al., 2020; Lorenzo-Dus and Bou-Franch, 2013). As Figure 1 illustrates, different from the carefully curated queries from traditional NLP benchmarks, real-world user-LLM queries tend to be less formal and frequently contain grammatical inconsistencies (Ouyang et al., 2023).
Failing to account for these linguistic variations risks excluding a broad segment of users from effective interaction, especially for users whose linguistic expressions fall outside the narrow patterns these systems are tuned on (Liang et al., 2023). Moreover, unlike standalone LLMs, RAG systems incorporate multiple interdependent components, making them susceptible to cascading errors arising at both the retrieval and generation stages (Asai et al., 2023; Yoran et al., 2024a; Kim et al., 2025). A truly robust RAG system should maintain consistent retrieval effectiveness and generation quality across the full spectrum of user linguistic variations.
@@
We present a large-scale systematic investigation of how variations in linguistic characteristics affect the robustness of RAG systems. We target diverse and prevalent variations commonly found in real-world user inputs that meaningfully challenge RAG systems (Park et al., 2024; Ouyang et al., 2023), namely, formality, readability, politeness, and grammatical correctness. These choices ensure our analysis covers stylistic, pragmatic, and structural aspects aligned with practical usage. By automatically rewriting queries across these dimensions, we analyze how linguistic variations impact each RAG system component, as well as potential cascading errors throughout the pipeline. Our evaluation encompasses two retrieval models, namely Contriever (Izacard et al., 2021) and ModernBERT (Nussbaum et al., 2024), and nine LLMs from three families (Llama 3.1, Qwen 2.5, Gemma 2) of varying scales across four open-domain Question Answering (QA) datasets: PopQA (Mallen et al., 2023), EntityQuestions (Sciavolino et al., 2022), MS MARCO (Bajaj et al., 2018), and Natural Questions (Kwiatkowski et al., 2019).
@@
Furthermore, RAG systems exhibit greater vulnerability to linguistic variations than LLM-only generations, suggesting cascading errors across components. The PopQA dataset shows a 22.53% average performance drop in RAG systems versus only 10.78% in LLM-only generations. These findings highlight the urgent need to improve the robustness of retrieval components when handling linguistically varied queries. We also find that while advanced RAG techniques, e.g., query expansion with HyDE (Gao et al., 2022) and documents reranking, tend to improve overall performance, they remain similarly vulnerable to linguistic variations.
@@ 2 Related Work
Prior research investigated retrieval system robustness to noisy queries (Campos et al., 2023; Chen et al., 2022, 2023b) and specific linguistic variations including word substitutions (Wu et al., 2022), aspect changes and paraphrasing (Penha et al., 2022), typographical errors (Zhuang and Zuccon, 2021), and grammatical variations (Long et al., 2024). Our work provides the first holistic evaluation of RAG systems’ robustness to diverse linguistic variations and uncovers cascading failures across the entire RAG pipeline.
@@
Prior research has examined impacts of syntactic perturbations Moradi and Samwald (2021); Singh et al. (2024), round-trip translation Bhandari and Chen (2023), politeness variations Yin et al. (2024), equivalent queries Cao et al. (2024) and scale of model size on robustness to grammatical errors (Yin et al., 2020; Hagen et al., 2024). Rawte et al. (2023) examined how formality and readability affect LLM performance in isolation. In contrast, our research investigates a broader spectrum of linguistic variations, generates data for each variation using LLM-based rewrites, and studies their compounding effects throughout the end-to-end RAG pipeline comprehensively.
@@
While RAG systems demonstrate impressive performance (Lewis et al., 2021; Gao et al., 2024) and reduced hallucinations (Mallen et al., 2023), vulnerabilities exist with increasing retrieval context noise (Chen et al., 2023a), irrelevant contexts (Yoran et al., 2024b), and noise impacts (Fang et al., 2024). Yang et al. (2025) analyzed how spurious features affect RAG perform. Cho et al. (2024) introduced document-level perturbations and evaluated RAG’s vulnerability to noisy documents. RbFT (Tu et al., 2025) introduce specialized training objectives to improve consistency under corrupted or adversarial retrieval contexts. These studies focus primarily on retrieved content noise rather than initial query; our work uniquely demonstrates how diverse linguistic variations in user queries compound throughout the RAG pipeline, exposing critical vulnerabilities in systems serving diverse users.
@@ 3 Robustness Evaluation Approach
Given a seed dataset $\mathcal{D}=\{(x_{1},y_{1}),(x_{2},y_{2}),\ldots\}$, where $x_{i}$ and $y_{i}$ indicate $i$-th input and output, respectively, we reformulate each query $x_{i}\rightarrow x_{i}^{\prime}$ based on four linguistic aspects. A robust RAG system, composed of a retriever $\mathcal{R}$ and a generator $\mathcal{G}$ operating on corpus $\mathcal{C}$, should maintain performance when processing linguistically varied inputs that are semantically equivalent, i.e., where the gold output remains unchanged semantically. For the retrieval component, we expect retrieved documents $\mathbf{D}_{i}=\mathcal{R}(x_{i},\mathcal{C})$ and $\mathbf{D}_{i}^{\prime}=\mathcal{R}(x_{i}^{\prime},\mathcal{C})$ to both contain the answer $y_{i}$. For generation, given retrieved documents $\mathbf{D}_{i},\mathbf{D}_{i}^{\prime}$, a robust system should produce: $\mathcal{G}(x_{i},\mathbf{D}_{i})\approx\mathcal{G}(x_{i}^{\prime},\mathbf{D}_{i}^{\prime})\approx y_{i}$.
@@ 3.1 Linguistic Variations
Formality in language lacks universal definition (Pavlick and Tetreault, 2016; Mosquera and Moreda, 2011; Fang and Cao, 2009), but encompasses situational factors (Hovy, 1987; Lahiri et al., 2011), grammar quality (Peterson et al., 2011), and specific linguistic elements like contractions (Heylighen and Dewaele, 1999). We quantify formality using the RoBERTa-based formality classifier from (Babakov et al., 2023).
@@
Readability quantifies text comprehensibility through linguistic complexity. We employ the widely-used Flesch Reading Ease Score (FRES; Flesch 1948) to assess readability (Rawte et al., 2023; Han et al., 2024), defined in A.1.
@@
In our work, we define grammatical correctness as the preservation of both grammaticality (Chomsky, 2002) and semantic fidelity. We alter the grammatical correctness through two approaches, inspired by Yin et al. (2020); Zhuang and Zuccon (2021); Lichtarge et al. (2019): (1) Typos, where random addition, deletion, or substitution operations are applied at a 20% probability per word, requiring an edit distance of at least 1; and (2) Round-trip translation (RTT) via English-Afrikaans-English using EasyNMT with opus-mt model, requiring the output to not be the exact same as the original query.
@@ 3.2 Query Rewriting
Grammar: RTT $\downarrow$ Original: Which company is HMS Blankney produced by?
@@
Grammar: Typos $\downarrow$ Original: when did the japanese river otter become extinct
@@
We systematically reformulate queries $x_{i}\rightarrow x_{i}^{\prime}$ across targeted linguistic dimensions while ensuring all rewrites satisfy dimension-specific thresholds and preserve semantic consistency (Appendix A.1).
Linguistic reformulation.
For each distinct dimension (formality, readability, and politeness), we sample $5,000$ original queries and use GPT-4o-mini (OpenAI et al., 2024) to generate rewrites using three different prompts (detailed in Appendix I). [^2] To verify that our experimental results are robust to rewriting models, we generate $500$ rewritten queries from the PopQA dataset using Llama-3.1-70B-Instruct. Shown in Appendix C, results confirm our findings with GPT-4o-mini. We design distinct query sets across all datasets to be less formal, less readable, or more polite—directions chosen because the well-curated original datasets contained predominantly formal, readable, and neutral language, allowing our modifications to explore more realistic linguistic variations. This creates $15,000$ rewritten samples per dimension for each dataset, over which we average results to account for prompt stochasticity. For grammatical correctness, we similarly use $5,000$ original queries but apply deterministic transformations through typo introduction and round-trip translation as described in Section 3.1, creating $5,000$ modified samples for each approach. Example rewrites are shown in Table 1.
Semantic consistency.
We define a rewritten query $x_{i}^{\prime}$ as semantically consistent with the original query $x_{i}$ if it preserves the original pragmatic intent, i.e., a human annotator would assign the same gold answer $y_{i}$ to both queries. This operational definition reflects the requirements that linguistic variation should not alter the underlying information need. To strictly enforce this, we apply a two-stage verification procedure. First, we automatically filter rewrites using a sentence-level semantic similarity threshold (>0.7>0.7) computed with MPNet-v2 (Song et al., 2020). Second, we conduct qualitative evaluation through manual annotation on $250$ queries across all linguistic variations, finding that 94.67% of the rewritten queries preserve the original answer. Further details of the annotation are provided in Appendix A.2.
For each linguistic dimension and dataset, we construct controlled comparative datasets $\mathcal{D}^{\prime}=\{(x_{1}^{\prime},y_{1}),(x_{2}^{\prime},y_{2}),\ldots\}$, enabling direct measurement of performance changes attributable solely to linguistic variation.
@@ 4.1 Benchmarks
In this work, we use four open-domain QA datasets as seed datasets $\mathcal{D}$ and evaluate the effects how linguistic variations of the original queries affect RAG systems. PopQA (Mallen et al., 2023) is a large-scale entity-centric open-domain QA dataset about entities with a wide variety of popularity. EntityQuestions (Sciavolino et al., 2022) is a set of simple, entity-rich questions based on facts of Wikidata. MS MARCO (Bajaj et al., 2018) contains questions derived from real user search queries from Bing’s search logs. Natural Questions (Kwiatkowski et al., 2019) contains questions consisting of real, anonymized, aggregated queries to the Google search engine. Both PopQA and EntityQuestions consist of well-structured, standardized, and simple queries, while queries from MS MARCO and Natural Questions exhibit free-form and arbitrary. We evaluate using PopQA’s test split, EntityQuestions’ dev split, and both dev and test splits from MS MARCO and Natural Questions. The retrieval is performed on the Wikipedia passage set used in DPR [^3] https://dl.fbaipublicfiles.com/dpr/wikipedia_split/psgs_w100.tsv.gz for the PopQA, EntityQuestions, and Natural Questions datasets, while the MS MARCO dataset uses its corresponding passage dataset (Bajaj et al., 2018).
@@ 4.2 Models
We evaluate the nine most advanced open-source instruction-tuned LLMs from three model families with various scales: Llama 3.1 (Grattafiori et al. 2024; 8 and 70 billion), Qwen 2.5Yang et al. 2024; Team 2024; 3, 7, 32, and 72 billion), and Gemma 2Team et al. 2024; 2, 9, and 27 billion). Detailed hyperparameter settings are included in the Appendix B.1.
We use few-shot prompting to ensure that the model outputs are in the correct format. For each dataset and linguistic characteristic, we include two question-answer pairs in the context: one random original query with its answer, and its corresponding linguistically rewritten version with the same answer. This balanced approach exposes the model to both original and rewritten query formats to ensure fairness. We also include the top five retrieved passages in the context. Detailed prompts are provided in Appendix J.
@@ 4.3 Metrics
The generation stage is assessed using a comprehensive set of metrics: Answer Match (AM) measures the percentage of the predictions that any substring of the prediction is an exact match of any of the ground truth answers, Exact Match (EM) measures the percentage of predictions that exactly match any of the ground truth answers, and F1 Score (F1) captures the harmonic mean of precision and recall in generated responses. We mainly report AM scores in the main paper because it better reflects answer correctness under linguistic and stylistic variaion. The full results could be found in Appendix F. As a complement to standard metrics, we also use LLM-as-a-Judge with GPT-5-mini (OpenAI, 2025).
@@ 5.1 Retrieval Analysis
Our analysis (Table 2) reveals significant linguistic fragility in retrieval systems, with performance degradation averaging 16.7% for Contriever and 13.3% for ModernBERT across all modifications. We hypothesize that ModernBERT’s increased robustness could largely be attributed to its diverse pre-training data, which forces the model to learn to process non-standard sequences and enhances its resilience to linguistic noise. Additionally, the masked language modeling objective utilized by ModernBERT could effectively train the model to denoise and reconstruct context from corrupted inputs. Results show highest sensitivity on PopQA (19.78% average impact), particularly to grammatical transformations (29.34% from RTT). MS MARCO exhibits the second-highest impact (16.78%), with striking sensitivity to formality changes (40.41% with Contriever), suggesting retrieval systems may be implicitly optimized for specific linguistic patterns, limiting effectiveness when handling diverse query variations.
@@
Retrieval performance drops independent of query length.
Our analysis, as shown in Table 2 demonstrates that query length changes do not directly correlate with retrieval performance. Queries with increased formality showed substantial length increases (+13.65 tokens) yet produced inconsistent performance impacts across datasets. Conversely, round-trip translated queries were marginally shorter (-0.26 tokens) but consistently caused significant performance degradation. This indicates retrieval models respond more to linguistic quality (grammatical correctness, readability) than to query length itself, highlighting the need for systems robust to linguistic variations rather than optimized for specific query lengths. We confirm this by analyzing semantic preservation using LLM-as-a-judge in Table 4 and a human evaluation of results in Table 5.
Scaling up number of documents improves performance.
[…]
Dataset Readability Formality Politeness Gram-RTT Gram-Typos
PopQA 19.41% 11.79% 6.75% 35.60% 21.20%
NQ 13.12% 4.27% 4.38% 22.74% 5.63%
MARCO 10.56% 4.05% 1.75% 15.21% 5.32%
Entity 13.20% 5.06% 2.86% 22.18% 10.49%
Table 4: RAG performance in LLM-as-a-Judge evaluation (ModernBERT + Qwen2.5-72B-Instruct), with relative percentage performance degradation. LLM-as-A-Judge evaluation shows a consistent degradation pattern.
5.2 Generation Analysis
The RAG experiment results on ModernBERT retriever and nine LLMs are presented in Table 3 with answer match (AM) scores. Table 4 shows the results with LLM-as-a-Judge evaluation. Overall, RAG systems show performance degradation on all linguistic variations.